OpenStudy (anonymous):

factorise xcu.-2xsq.-x+2

OpenStudy (anonymous):

\[x ^{3} - 2x ^{2}-x+2 =0\] \[x ^{2}(x-2) - (x-2) = 0\] \[(x-2)(x ^{2}-1)=0\] (x-2)(x-1)(x+1)=0

OpenStudy (anonymous):

thank you